OpenStudy (ccchristian1):

write the equation of a line that is perpendicular to the given line and that passes through the given point. y-2=7/3(x+5);(-4,9)

OpenStudy (ccchristian1):

a.y-9=-3/7(x-4) b.y-9=3/7(x-4) c.y-9=-3/7(x+4) d.y-9=3/7(x+4)

OpenStudy (ccchristian1):

@ivettef365

OpenStudy (ivettef365):

ok, what is the slope on the actual equation?

OpenStudy (ivettef365):

do you know?

OpenStudy (ccchristian1):

7/3

OpenStudy (ivettef365):

y - y1 = m (x - x1) slope is the m position you have y - 2 = 7/3 (x + 5)

OpenStudy (ccchristian1):

am i right with slope ?

OpenStudy (ivettef365):

ok good, so to be perpendicular what is the opposite reciprocal

OpenStudy (ivettef365):

yes you are right 7/3

OpenStudy (ccchristian1):

so its c?

OpenStudy (ivettef365):

yes

OpenStudy (ccchristian1):

using the stuff u taught me.. i got c

OpenStudy (ccchristian1):

alright awesome!

OpenStudy (ccchristian1):

thanks